An airplane flew for 4 hours with a 20 km/h tail wind. The return flight against the same wind took 5 hours. Find the speed of the plane in still air.
Answer by Paul(988)

lET THE SPEED OF THE PLANE BE X
Equation:
4(x+20)=5(x-20)
4x+80=5x-100
-x=-180
x=180
Hence, the speed of the plane is 180mph.
Paul.